The sum of two numbers is 82. Their difference is 24. Write a system of equations that describes this situation. Solve by elimination to find the two numbers.
Found 2 solutions by 303795, checkley71:
Answer by 303795(562)

Let one number be x and the second number y
x + y = 82
x - y = 24
To solve by elimination add the two equations together to get
2x = 106
x = 53
If one number is 53 substitute x=53 into the first equation.
53 + y = 82
y=82-53
y=29
Check in the other equation
53-29 = 24.

Answer by checkley71(8403)

X+Y=82
X-Y=24 NOW ADD THESE 2 EQUATIONS
--------------------------------------------
2X=106
X=106/2
X=53 SOLUTION
53+Y=82
Y=82-53
Y=29 SOLUTION
PROOF
53-29=24
24=24